Hilda Ltd uses the standard costing system. The standards for Product Z100 are as follows:

Material

4 kilograms at R2.50 per kilogram

Labour

4 hours at R100 per hour

Variable overheads

R40 per labour hour

Fixed overheads

R48 000

Normal production

24 000 units per month

Actual production for the month of November 2016 on Product Z100 is:

Material

81 000 kilograms used at R2.50 per kg

Labour

78 000 hours worked at R96 per hour

Variable overheads

R44 per labour hour

Fixed overheads

R53 000

Normal production

20 000 units

Required: 1.1 Calculate and comment on the following variances:

1.1.1 Raw material usage variance. (6)

1.1.2 Fixed overhead spending variance. (6)

1.1.3 Variable overhead efficiency variance. (6)

1.2 Explain the underlying causes for variances. (6)

1.3 Discuss the merits of Budgeting. (6)
Answer #1
1.1.1 Raw material usage variance =
(Actual usage - Standard usage ) * Standard cost per unit
From the given info,
Standard cost per unit = R 2.5
Standard usage = 20000 units * 4 kg = 80000 kgs
Actual usage = 81000 kgs
Raw material usage variance = (81000 - 80000) * 2.5 = 2500 (U)
1.1.2 Fixed overhead spending variance =
Actual fixed overheads - Budgeting fixed overheads
From the given info,
Actual fixed overheads = R 53000
Budgeted fixed overheads = R 48000
Fixed overhead spending variance = 53000 - 48000 = 5000 (U)
1.1.3 Variable overhead efficiency variance =
Number of hours of variable production overhead per unit - Number of hours budgeted
Budgeted variable overheads = R 40 per labour hour
Budgeted labour hours per unit = 4 hours
Budgeted labour hours for actual production of 20000 units = 4 * 20000 = 80000 hours
Actual labour hours = 78000 hours
Variable overhead efficiency variance = 78000 - 80000 = 2000 (F)

1.1.4 Causes of variances:

1 Lack of due care in usage of materials
2 Abnormal wastage through pilferage, etc.
3 Inefficiency in production due to improper method or lack of necessary skill in workmen
4 Use of a material-mix other than the standard mix
5 Purchase of inferior materials or change in quality of materials
6 Yield from materials in case excess of or less than that provided as the standard yield

1.1.5 Advantages of budgeting:

1 Budgetary control aims at maximisation of profits through effective planning and control of income and expenditure - directing capital and resources to the best and most profitable channel
2 It provides a clear definition of the objective and policies of the concern and a tool for objecting these policies to periodic examination.
3 There is a planned approach to expenditure and financing of the business so that economy is affected in the utilisation of funds to the optimum benefit of the concern.
4 The method of evaluating performance against budgets provides a suitable basis for establishing incentive system of remuneration by results as also spotting people with exceptional qualities of leadership and management
5 Budgets are the fore-runners of standard costs in the sense that they create necessary conditions to suit setting up of standard costs.
6 It cultivates in the management the habit of thinking ahead - making careful study of the problems in advance before taking decisions.